# ORFS-agent: Tool-Using Agents for Chip Design Optimization

### v1

@@ Abstract
Machine learning has been widely used to optimize complex engineering workflows across numerous domains. In the context of integrated circuit design, modern flows (e.g., going from a register-transfer level netlist to physical layouts) involve extensive configuration via thousands of parameters, and small changes to these parameters can have large downstream impacts on desired outcomes - namely design performance, power, and area. Recent advances in Large Language Models (LLMs) offer new opportunities for learning and reasoning within such high-dimensional optimization tasks. In this work, we introduce ORFS-agent, an LLM-based iterative optimization agent that automates parameter tuning in an open-source hardware design flow. ORFS-agent adaptively explores parameter configurations, demonstrating clear improvements over standard Bayesian optimization approaches in terms of resource efficiency and final design metrics. Our empirical evaluations on two different technology nodes and a range of circuit benchmarks indicate that ORFS-agent can improve both routed wirelength and effective clock period by over 13%, all while using 40% fewer optimization iterations. Moreover, by following natural language objectives to trade off certain metrics for others, ORFS-agent demonstrates a flexible and interpretable framework for multi-objective optimization. Crucially, ORFS-agent is modular and model-agnostic, and can be plugged in to any frontier LLM without any further fine-tuning.
@@ I Introduction
Large Language Models (LLMs) have reshaped AI, excelling at natural-language generation, question answering, and zero-/few-shot learning [7]. Their emergent reasoning spans mathematical problem solving, code generation, and multi-step orchestration via agent frameworks [12]. LLMs integrate into agent systems [50], pairing API calls, predefined functions (“tool use”), and context windows to meet language-defined goals. These agents already serve fields from healthcare to voice-controlled computing.
Research on LLMs has seen major advances in neural architecture search [13] [32], combinatorial optimization [5] [37], and heuristic selection [6] [28], all relevant to chip design sub-tasks. Chip design relies on a complex EDA flow that turns high-level hardware descriptions into manufacturable layouts. Each stage exposes hundreds or thousands of hyperparameters (timing constraints, placement strategies, routing heuristics, etc.) whose joint settings strongly affect power, performance, and area (PPA). At advanced nodes, design complexity and the parameter space balloon further [1].
The Concrete Problem. We target LLM-driven black-box optimization across varied objectives, circuits, and technology nodes. In ML terms, this is context-based Bayesian optimization: LLM agents exploit metadata (e.g., log files) as heuristics. Traditional Bayesian optimization (BO) tunes EDA parameters well, yet pure black-box methods miss domain knowledge and need explicit objectives. LLMs, with broad “world knowledge,” use chain-of-thought to iteratively propose and refine parameters [62] [53].
LLM agents interface with chip-design tools, parse intermediate results, call domain APIs, and adjust flow parameters in a human-like loop. They also slot into existing heuristic frameworks (e.g., BO) to boost efficiency.
@@
Alternative approaches avoid retraining, and instead wrap or retrieve around frozen SOTA models. EDA-Copilot [4], Ask-EDA [47], and ORAssistant [25] combine retrieval-augmented generation with off-the-shelf chat models to answer tool-usage questions, while surveys (LLM4EDA [62], “Future or Mirage?” [60], and the taxonomy in [45]) highlight the growing appeal of lightweight, model-agnostic agents.
@@
Our Contributions. We introduce ORFS-agent, an LLM-driven autotuner for an open-source chip-design flow that tops Bayesian optimization (BO) with fewer flow runs. Our main contributions are:
  • We embed ORFS-agent in OpenROAD-flow-scripts (ORFS)[70]. It schedules parallel ORFS runs, reads intermediate metrics (wirelength, timing, power), and iteratively refines parameters to optimize post-route results. Leveraging ORFS’s METRICS2.1 logging[22], the agent skips bad configurations and gives robust, efficient chip design optimization. Modularity allows instantly switching models.
  • ORFS-agent obeys natural-language instructions to weight or constrain design metrics (e.g., favor timing over wirelength) and taps existing BO libraries inside its tool-use loop for extra lift.
  • On ASAP7 and SKY130HD (three circuits each), ORFS-agent trims ORFS calls and improves wirelength and effective clock period by $\geq 13\%$ over baseline, beating the BO-based ORFS AutoTuner [22].
@@
In the following, Section II reviews BO and recent LLM-EDA work. Section III gives details of ORFS-agent—prompts, data, iterations. Section IV gives experiments. Section V concludes and sketches future work. Anonymized code is open-sourced for inspection at [71].
@@ II Related Work
We organize our discussion of related work into three subsections, covering Open-Source EDA, Bayesian Optimization, and LLMs for Optimization in EDA in that order. Additional discussion of EDA flow and OpenROAD is given in Appendix A.
@@ II-A Open-Source EDA and OpenROAD as an ML Testbed
We target LLM/ML EDA research based on the permissively open-sourced OpenROAD[3], where code, logs and data can be made public to enable reproducible training and evaluation. Tcl/Python APIs expose every flow stage, for capture of live metrics and debug – vital for ML agents. With its RTL-to-GDSII scope, transparency and active community, OpenROAD provides an excellent testbed. ORFS wraps OpenROAD in editable JSON/Tcl for large-scale experiments[70]. Other tools build on OpenROAD, e.g., DREAMPlace yields 30×30$\times$30 × faster placement than OpenROAD’s RePlAce via CUDA acceleration [30].
@@ II-B Bayesian Optimization in Flow Tuning
Bayesian optimization (BO) is a widely used approach for parameter autotuning in EDA flows [22] including in commercial tools such as Synopsys DSO.ai [68] and Cadence Cerebrus [65]. It enjoys efficiency in terms of samples and solution space search, but faces drawbacks:
  • Contextual and Knowledge Limitations: BO frameworks typically do not incorporate domain knowledge unless manually encoded (e.g., at PDK level). Further, they do not adapt from past results.
  • Explicit Objective Functions: BO frameworks require explicitly defined objective functions, possibly missing complex or subjective criteria. This can hinder effectively balancing PPA in chip design.
  • Scalability Challenges: High-dimensional spaces and large designs can combinatorially overwhelm traditional BO approaches [57].
We seek flexible, context-aware and data-driven approaches to parameter tuning. Large Language Models offer a promising solution.
II-C Large Language Models (LLMs) in EDA Optimization
@@
Contextual knowledge and adaptivity. Unlike black-box Bayesian optimization (BO), LLMs ingest circuit data, retrieve literature, and adjust strategies on the fly. They read partial metrics (e.g., post- CTS timing) and learn user preferences without explicit objectives [40].
Recent progress. Advisor systems LLM4EDA and ChatEDA guide placement and timing fixes [62] [53]. Startups Silogy, Silimate, Atopile, Diode Computers and ChipAgents extend AI across design tasks. New methods such as LLAMBO [31, 69], forum posts on r/MachineLearning, and the LAD Symposium reinforce the trend. Multi-fidelity BO and meta-surrogates [14, 26, 46] improve sampling efficiency yet still treat flows as opaque; by contrast, LLMs promise context-aware tuning. Large-scale industrial adoption of LLMs for EDA can be seen, e.g., NVIDIA’s MARCO framework shows graph-based multi-agent task solving for design flows[66]. Open problems in LLM-based hardware verification are mapped in [54].
@@ II-D Working with LLMs and Agent Frameworks
LLMs need scaffolding beyond raw prompts. Key techniques are:
@@
  • Tool-use / agent frameworks [44, 15, 43] let an LLM call shells or solvers, parse error logs [20], and orchestrate full flows [27].
@@
More modern advances. OPRO treats the model itself as the optimizer [58]. LLAMBO warm-starts BO with zero-shot hints [31]. ADO-LLM fuses analog priors with BO loops [59]. ChatEDA and the newer EDAid system wire agents straight into complete RTL-to-GDSII flows [53] [55]. RL efforts such as AutoDMP and Google’s graph-placement policy show tractability of end-to-end layout [2] [10] [38].
Tool-centric agent stacks. Browser agents such as WebArena surf, click, and scrape constraints for downstream solvers [61]. Terminal aides—ShellGPT and Devin—run make, git, and yosys on command, closing the loop between code and silicon [73] [67]. API routers like ToolLLM, Gorilla, and OpenAgents map natural-language tasks onto thousands of REST calls and plugins [44] [43] [56]. AutoGPT-style self-looping controllers review results and re-issue actions [48], while the AI-Scientist pipeline shows how the same pattern can automate hypothesis, experiment, and paper generation [74]. At code level, AlphaCode 2 beats 85% of Codeforces competitors [64]; RL-trained VeriSeek and retrieval-guided logic synthesis agents trim Verilog area-delay in minutes [35] [8]. Together, browser, tool, and terminal hooks turn LLMs into end-to-end optimization engines.
@@ III Proposed Methodology
We now introduce the integration framework of our ORFS-agent with the OpenROAD-flow-scripts (ORFS). We then present an explanation of a single iteration of the ORFS-agent loop.
@@ III-A Overview of ORFS-agent
where each $f_{i}$ is a design metric (e.g., routed wirelength, worst negative slack) and each $\alpha_{i}$ is a nonnegative real weight. This formulation allows optimization of either a single metric or a weighted combination of multiple metrics. Then, ORFS-agent iteratively modifies the configuration file (config.mk) for ORFS and SDC files to optimize the specified objectives. The ORFS configuration file defines adjustable parameters (knobs) for placement, clock tree synthesis and routing. The SDC file specifies timing constraints for the target design, such as clock period. The outputs of ORFS-agent are the optimized configuration file for ORFS and the optimized SDC file; together these optimize the user-specified objectives under given constraints.
ORFS-agent relies on an LLM (e.g., GPT, Claude, Llama) that can: (i) read and modify files and logs, (ii) optionally invoke external tools via function calling, i.e., specify argument values for pre-existing functions (e.g., Python functions), and (iii) propose new parameter values to improve the objective under given constraints. There is a key distinction between direct LLM use and the use of LLM-based agents with function-calling capabilities. In the function-calling paradigm, the LLM is given a goal (e.g., “write a function to compute a matrix eigendecomposition”) along with a list of tools, which are black-box functions the LLM can invoke to facilitate task completion. For example, a tool named ${\mathbf{ISDIAGONALIZABLE}}$ determines whether a matrix $M$ is diagonalizable and returns a Boolean value. The tools used in ORFS-agent fall into the following three categories:
@@ III-B Overall Iteration Structure with Example and Tool Usage
We maintain a GLOBALCONTEXT that collects relevant parameters and partial metrics across steps. The initialization of this GLOBALCONTEXT at the first iteration is also the LLM prompt; this initialization, along with procedures for update between iterations for GLOBALCONTEXT, is separately discussed in Appendix E. Let K denote the number of parallel runs per iteration, and let TIMEOUT specify the maximum runtime for each batch. Below, we show an example scenario with $\texttt{K}=25$, $\texttt{TIMEOUT}=30$ and parameters (Core Util, Clock Period) can be integrated, along with possible function-calling tools. Let us consider the $j$-th iteration. Assume that we target the routed wirelength and take the CTS wirelength as a surrogate (recalling that the CTS stage occurs earlier and thus will likely not time out).
@@
  • Example: For $\texttt{K}=25$ and $\texttt{TIMEOUT}=30$ minutes, we launch 25 parallel runs, each exploring a different (Core Util, Clock Period) setting. All will run for at most $30$ minutes, producing log files.
@@
Stage 2: READ. Gather JSON logs from the files generated by all RUN steps up to this iteration, which may contain partial surrogate metrics, e.g., from CTS stages as well as ground truth metrics from finished runs.
  • Example: The agent collects $25\times j$ JSON files, with the routed wirelength (ground truth) or the earlier, but rough CTS wirelength (surrogate).
@@
Stage 3: COLLATE. Merge the logs into a single dataset, storing outcomes for all K ×$\times j$ runs (e.g., a table or array).
@@
  • Tool use: The agent calls pre-defined INSPECT tools to examine distributions, which allows sidestepping inspecting the values in context. For example, $\mathbf{InspectDistribution(ARR,X,Y,Y_{surrogate})}$ outputs the distributional relation between an input $\mathbf{X}$ and targets $\mathbf{Y}$ with surrogate targets $\mathbf{Y_{surrogate}}$. Here we have Clock Period as $\mathbf{X}$, CTS and routed wirelength as $\mathbf{Y_{surrogate},Y}$ and $\mathbf{ARR}$ from the COLLATE step.
@@
  • Example usage: The agent calls a OPTIMIZE tool, e.g., a Gaussian Process ($\mathbf{GP}$) with the COLLATE-level data, to get $200$ values of candidate Clock Period, Core Util pairs, using a Matern kernel, and Expected Improvement as search strategy.
@@
Stage 6: AGGLOMERATE (optional). If more than K candidates are generated in the OPTIMIZE phase, then reduce to exactly K, often by maximizing diversity or coverage.
  • Example: Since the $\mathbf{GP}$ produces $200$ new (Core Util, Clock Period) pairs; we reduce to 25 via selecting through a valid AGGLOM tool, $\mathbf{EntropySelect(ARR,quality,points)}$. This function returns $\mathbf{points}$, i.e., $25$ pairs of (Core Util, Clock Period) that it thinks are optimal as a set w.r.t. the provided $\mathbf{quality}$ metric (wirelength improvement) on the same $\mathbf{ARR}$ from COLLATE.
@@
  • Example: Generate 25 updated config/SDC files (or commands), each specifying one of the chosen (Core Util, Clock Period) combinations.
@@
This process repeats until the stopping criteria—generally, total serial iteration count—are met. The outputs of stages READ to AGGLOMERATE enter the context window of the ORFS-agent in each iteration, but not necessarily into GLOBALCONTEXT. Context management and tool-level details appear in Appendix E.

### v2

@@ Abstract
Machine learning has been widely used to optimize complex engineering workflows across numerous domains. In the context of integrated circuit design, modern flows (e.g., going from a register-transfer level netlist to physical layouts) involve extensive configuration via thousands of parameters, and small changes to these parameters can have large downstream impacts on desired outcomes – namely design performance, power, and area. Recent advances in Large Language Models (LLMs) offer new opportunities for learning and reasoning within such high-dimensional optimization tasks. In this work, we introduce ORFS-agent, an LLM-based iterative optimization agent that automates parameter tuning in an open-source hardware design flow. ORFS-agent adaptively explores parameter configurations, demonstrating clear improvements over standard Bayesian optimization approaches in terms of resource efficiency and final design metrics. Our empirical evaluations on two different technology nodes and a range of circuit benchmarks indicate that ORFS-agent can improve both routed wirelength and effective clock period by over 13%, all while using 40% fewer optimization iterations. Moreover, by following natural language objectives to trade off certain metrics for others, ORFS-agent demonstrates a flexible and interpretable framework for multi-objective optimization. Crucially, ORFS-agent is modular and model-agnostic, and can be plugged in to any frontier LLM without any further fine-tuning.
@@ I Introduction
Large Language Models (LLMs) have reshaped AI, excelling at natural-language generation, question answering, and zero-/few-shot learning [7]. Their emergent reasoning spans mathematical problem solving, code generation, and multi-step orchestration via agent frameworks [12]. LLMs integrate into agent systems [53], combining API calls, predefined functions (“tool use”), and context windows to meet language-defined goals. Agents already serve a range of fields from healthcare to voice-controlled computing.
Research on LLMs has seen advances in neural architecture search [13] [33], combinatorial optimization [5] [38], and heuristic selection [6] [29], all of which are relevant to chip design tasks. Chip design relies on a complex EDA flow that turns high-level hardware descriptions into manufacturable layouts. Each stage exposes hundreds or thousands of hyperparameters (timing constraints, placement strategies, routing heuristics, etc.) whose joint settings strongly affect power, performance, and area (PPA). At advanced nodes, design complexity and the design flow parameter space balloon further [1].
The concrete problem. We target LLM-driven black-box optimization across varied objectives, circuits, and technology nodes. In ML terms, this is context-based Bayesian optimization: LLM agents exploit metadata (e.g., log files) as heuristics. Traditional Bayesian optimization (BO) can tune EDA parameters well, but pure black-box methods miss domain knowledge and need explicit objectives. By contrast, LLMs have broad “world knowledge” and can use chain-of-thought to iteratively propose and refine parameters [65] [56].
LLM agents interface with chip design tools, parse intermediate results, call domain APIs, and adjust flow parameters in a human-like loop. They can boost the efficiency of heuristic frameworks (e.g., BO).
@@
Other methods avoid retraining, and instead wrap or retrieve around frozen SOTA models. EDA-Copilot [4], Ask-EDA [50], OpenROAD-Assistant [48], RAG for EDA-QA [45] and ORAssistant [26] combine retrieval-augmented generation with off-the-shelf models for tool-usage questions, while surveys  [65] [63] [47] highlight the draw of lightweight, model-agnostic agents.
@@
Our contributions. We develop ORFS-agent, an LLM-driven autotuner for an open-source chip-design flow that can surpass Bayesian optimization using fewer flow runs. Our main contributions are:
  • We embed ORFS-agent in OpenROAD-flow-scripts (ORFS)[79]. It schedules parallel ORFS runs, reads intermediate metrics (wirelength, timing, power), and iteratively refines parameters to optimize post-route results. Leveraging ORFS’s METRICS2.1 logging[22], the agent skips bad configurations and gives robust, efficient chip design optimization. Modularity allows frictionless switching of models.
  • ORFS-agent obeys natural-language instructions to weight or constrain design metrics (e.g., to favor timing over wirelength) and is able to leverage existing BO libraries inside its tool-use loop.
[…]
@@
In the following, Section II reviews BO and recent LLM-EDA work. Section III gives details of ORFS-agent – prompts, data, and iterations. Section IV gives experiments. Section V concludes and sketches future work. Our code is open-sourced at [80].
@@ II Related Work
We taxonomize related work into three area: Open-Source EDA, Bayesian Optimization, and LLMs for Optimization in EDA. Appendix A gives additional discussion of EDA flow and OpenROAD.
@@ II-A Open-Source EDA and OpenROAD as an ML Testbed
We target LLM/ML EDA research based on the permissively open-sourced OpenROAD[3], whose code, logs and data can be made public to enable reproducible training and evaluation. Tcl/Python APIs expose every flow stage for capture of live metrics and debug, which is vital for ML agents. With its RTL-to-GDSII scope, transparency and active community, OpenROAD provides a robust testbed. ORFS wraps OpenROAD in editable JSON/Tcl for large-scale experiments[79].
@@ II-B Bayesian Optimization in Flow Tuning
Bayesian optimization (BO) is a widely used approach for parameter autotuning in EDA flows [22], including in commercial tools such as Synopsys DSO.ai [75] and Cadence Cerebrus [70]. BO is efficient in terms of samples and solution space search, but has several drawbacks.
  • Contextual and knowledge limitations. BO frameworks typically do not incorporate domain knowledge unless manually encoded (e.g., at PDK level). Further, they do not adapt from past results.
  • Explicit objective functions. BO frameworks require explicitly defined objective functions, possibly missing complex or subjective criteria. This can hinder effectively balancing PPA in chip design.
  • Scalability challenges. High-dimensional spaces and large designs can combinatorially overwhelm traditional BO approaches [60].
We seek flexible, context-aware and data-driven approaches to parameter tuning. Large Language Models offer a promising foundation.
II-C Large Language Models in EDA Optimization
@@
Contextual knowledge and adaptivity. Unlike black-box Bayesian optimization, LLMs ingest circuit data, retrieve literature, and adjust strategies on the fly. They read partial metrics (e.g., post-clock tree synthesis (-CTS) timing) and learn user preferences without explicit objectives [41].
Recent progress. Advisor systems LLM4EDA [65] and ChatEDA [56] guide placement and timing fixes. Startups Silogy [85], Silimate [84], Atopile [69], Diode Computers [74] and ChipAgents [71] extend AI across design tasks. New methods such as LLAMBO [32] [77] and works presented at ICLAD [76] and MLCAD [78] Symposia reinforce the trend. Multi-fidelity BO and meta-surrogates [14] [27] [49] improve sampling efficiency yet still treat flows as opaque; by contrast, LLMs promise context-aware tuning. Large-scale industrial adoption of LLMs for EDA can be seen, e.g., NVIDIA’s MARCO framework shows graph-based multi-agent task solving for design flows[72]. Open problems in LLM-based hardware verification are mapped in [57].
@@ II-D Working with LLMs and Agent Frameworks
LLMs require scaffolding beyond raw prompts. Techniques include:
@@
  • Tool-use/agent frameworks [46] [15] [44] let an LLM call shells or solvers, parse error logs [20], and orchestrate full flows [28].
@@
More modern advances. OPRO [61] treats the model itself as the optimizer. LLAMBO warm-starts BO with zero-shot hints [32]. ADO-LLM fuses analog priors with BO loops [62]. ChatEDA [56] and the newer EDAid [58] systems wire agents straight into complete RTL-to-GDSII flows. RL efforts such as AutoDMP [2] [10] and Google’s graph-placement policy [39] show tractability of end-to-end layout.
Tool-centric agent stacks. Browser agents such as WebArena [64] surf, click, and scrape constraints for downstream solvers. Terminal aides such as ShellGPT [82] and Devin [73] run make, git, and yosys on command, closing the loop between code and silicon. API routers such as ToolLLM [46], Gorilla [44], and OpenAgents [59] map natural-language tasks onto thousands of REST calls and plugins. AutoGPT [51]-style self-looping controllers review results and re-issue actions, while the AI-Scientist [83] pipeline shows how the same pattern can automate hypothesis, experiment, and paper generation. At code level, AlphaCode 2 [66] beats 85% of Codeforces competitors; RL-trained VeriSeek [36] and retrieval-guided logic synthesis agents [8] trim Verilog area-delay in minutes. Together, browser, tool, and terminal hooks turn LLMs into end-to-end optimization engines.
Our work explores how these engines might advance the state of the art in heuristics-driven applied optimization domains such as EDA.
@@ III Proposed Methodology
We now introduce the integration framework of our ORFS-agent with OpenROAD-flow-scripts (ORFS). We then present an explanation of a single iteration of the ORFS-agent loop.
@@ III-A Overview of ORFS-agent
where each $f_{i}$ is a design metric (e.g., routed wirelength, worst negative slack) and each $\alpha_{i}$ is a nonnegative real weight. This formulation allows optimization of either a single metric or a weighted combination of multiple metrics. Then, ORFS-agent iteratively modifies the configuration file (config.mk) for ORFS and SDC files to optimize the specified objectives. The ORFS configuration file defines adjustable parameters (knobs) for floorplanning, placement, clock tree synthesis and routing. The SDC file specifies timing constraints for the target design, such as clock period. The outputs of ORFS-agent are the optimized configuration file for ORFS and the optimized SDC file; together, these optimize the user-specified objectives under given constraints.
ORFS-agent relies on an LLM (e.g., GPT, Claude, Llama) that can (i) read and modify files and logs; (ii) optionally invoke external tools via function calling, i.e., specify argument values for pre-existing functions (e.g., Python functions); and (iii) propose new parameter values to improve the objective under given constraints. There is a key distinction between direct LLM use and the use of LLM-based agents with function-calling capabilities. In the function-calling paradigm, the LLM is given a goal (e.g., “write a function to compute a matrix eigendecomposition”) along with a list of tools, which are black-box functions the LLM can invoke to facilitate task completion. For example, a tool named ${\mathbf{ISDIAGONALIZABLE}}$ determines whether a matrix $M$ is diagonalizable and returns a Boolean value. The tools used in ORFS-agent fall into the following three categories.
@@ III-B Overall Iteration Structure with Example and Tool Usage
[…]
@@
[…]
@@
Stage 2: READ. Gather JSON logs from the files generated by all RUN steps up to this iteration. These may contain partial surrogate metrics, e.g., from the CTS stage, or ground truth metrics from finished runs.
  • Example: The agent collects $25\times j$ JSON files, with routed wirelengths (ground truth) or earlier rough CTS wirelengths (surrogate).
@@
Stage 3: COLLATE. Merge the logs into a single dataset, storing outcomes for all K $\times$ $j$ runs (e.g., a table or array).
@@
  • Tool use: The agent calls pre-defined INSPECT tools to examine distributions; this allows sidestepping of in-context inspection of values. For example, $\mathbf{InspectDistribution(ARR,X,Y,Y_{surrogate})}$ outputs the distributional relation between an input $\mathbf{X}$ and targets $\mathbf{Y}$ with surrogate targets $\mathbf{Y_{surrogate}}$. Here we have Clock Period as $\mathbf{X}$, CTS and routed wirelength as $\mathbf{Y_{surrogate},Y}$ and $\mathbf{ARR}$ from the COLLATE step.
@@
  • Example: The agent calls an OPTIMIZE tool, e.g., a Gaussian process ($\mathbf{GP}$) with the COLLATE-level data, to get 200200200 values of candidate Clock Period, Core Util pairs, using a Matérn kernel, and Expected Improvement as search strategy.
@@
Stage 6: AGGLOMERATE (optional). If more than K candidates are generated in the OPTIMIZE phase, then reduce to exactly K, e.g., by maximizing diversity or coverage.
  • Example: Since the $\mathbf{GP}$ produces 200200200 new (Core Util, Clock Period) pairs; we sub-select to 25 via a valid AGGLOM tool, $\mathbf{EntropySelect(ARR,quality,points)}$. This function returns $\mathbf{points}$, i.e., 252525 pairs of (Core Util, Clock Period) that are considered set-level optimal with respect to the provided $\mathbf{quality}$ metric (wirelength improvement) on the same $\mathbf{ARR}$ from COLLATE.
@@
  • Example: Generate 25 updated config/SDC files or commands, each having one of the chosen (Core Util, Clock Period) combinations.
@@
This process repeats until the stopping criteria – generally, total serial iteration count – are met. The outputs of stages READ through AGGLOMERATE enter the context window of ORFS-agent in each iteration, but do not necessarily enter the GLOBALCONTEXT. Context management and tool-level details are given in Appendix E.